# Subject-Oriented Video Captioning

### v1

@@ 5.1 Baseline Methods
5.2 LSTM-based Subject-Oriented Methods
The structure of the LSTM-based method is shown in Figure 6 (a). , HMN and GL-RG belong to this category. They both consist of a Video Encoder and an LSTM Decoder to process and generate the video captions, respectively. The caption generation process is as follows:
[…]
$F_{v}=E_{v}(V).$ (1)
The structure of the encoder varies across methods. For instance, HMN employs a complex encoder comprising both Transformer and BiLSTM. For ease of representation, we uniformly denote as $E_{v}$.
Then, video features $F_{v}$ and the word $x_{t-1}$ generated at the previous time step are fed into an ${{\rm LSTM}}$ to predict word $x_{t}$ for the current time step:
${}x_{t}={{\rm LSTM}}(F_{v},x_{t-1}).$ (2)
[…]
$F_{s}=E_{s}(I).$ (3)
The encoder’s structure includes a pre-trained feature extractor and a linear layer for dimensionality reduction. For ease of description, we uniformly denote this entire structure as $E_{s}$. Then, we concatenate the encoded subject features with the original encoded video features
$F_{m}=Concat(F_{v},F_{s}),$ (4)
which is fed into the LSTM decoder Eq. 2 for the next word prediction.
@@ 5.2 LSTM-based Subject-Oriented Methods
AR-B [28]_ori_perf 2021 48.7 33.3 - 91.8 40.5 28.7 - 49.1
AR-B 2023 27.4 22.4 50.5 69.7 17.0 16.6 38.7 43.3
AR-B_ext 2023 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN [32]_ori_perf 2022 59.2 37.7 75.1 104.0 43.5 29.0 62.7 51.5
HMN 2023 29.8 22.9 51.2 77.7 17.1 16.5 39.1 44.9
HMN_ext 2023 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG [27]_ori_perf 2022 45.5 30.1 62.6 94.3 55.5 37.8 74.7 51.2
GL-RG 2023 34.5 26.1 54.4 86.7 18.3 17.3 39.1 43.8
GL-RG_ext 2023 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert [9]_ori_perf 2022 55.7 39.7 75.7 109.4 41.9 29.9 62.1 53.8
SwinBert 2023 33.4 25.7 54.2 89.6 18.2 17.4 40.4 46.9
SwinBert_ext 2023 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 2: Results on our Subject-Oriented MSVD and MSR-VTT datasets. Rows with grey background denote results on original MSVD and MSRVTT datasets. The best results are highlighted in bold.
5.3 Transformer-based Subject-Oriented Methods
The video captioning methods based on the Transformer architecture can be further divided into two categories: Transformer encoder-based methods and Transformer decoder-based methods. Specifically, for the first category of methods, the model’s decoder primarily utilizes the architecture of the ${{\rm TransEncoder}}$. As shown in Figure 6 (b), text tokens $T$ and video tokens $F_{v}$ are concatenated along the sequence dimension. Each of these tokens serves as the query, key, and value, which are fed into the transformer encoder to generate target text:
$T_{gen}=\{x_{t}\}_{t=1}^{N}={{\rm TransEncoder}}(Concat(T,F_{v})),$ (5)
$T_{gen}$ denotes the generated caption sequence of length $N$. For the second category of methods, the model’s decoder mainly adopts the structure of the ${{\rm TransDecoder}}$. Text tokens $T$ serves as the query, with video tokens $F_{v}$ acting as keys and values used for generating text:
$T_{gen}=\{x_{t}\}_{t=1}^{N}={{\rm TransDecoder}}(T,F_{v},F_{v}).$ (6)
For the above transformer-based methods, we concatenate subject tokens with video tokens along the sequence dimension. These models can utilize their attention mechanism to automatically learn representations to facilitate the generation of desired captions.
6 Experiments
6.1 Implementation Details
We adopt the official codes of the four state-of-the-art methods and re-train them on our dataset using default hyperparameters considering they have similar data sizes. As for the object encoder employed in our new approach, we experimented with both Faster-RCNN [18] and CLIP [17]. The size of object features is 2048 obtained by Faster-RCNN. In the case of the pre-trained CLIP model, we fed the cropped regions directly into the image encoder with a VIT-L/14 backbone, leading to a feature dimension of 768.
6.2 Quantitative Results
Baselines Performance
In Table 2, we present the performance of the four baseline models and our extended models based on them. For a clear comparison, we also included their performance on the original MSVD and MSRVTT datasets, which are presented in rows with grey backgrounds.
The results show that: (I) The four state-of-the-art methods encounter a large performance drop on our new task compared to the performance on the original video captioning task. For example, on our subject-oriented MSVD, the CIDEr score of AR-B decreases by 31.7% relative to itself, while HMN reduces by 33.8% compared to its own performance. These indicate our subject-oriented video captioning task poses new challenges to existing methods. To succeed, they should have the ability to be aware of the specified target, and the ability to distinguish related activities. (II) Our extension of the four state-of-the-art methods brings large improvement. For example, on subject-oriented MSRVTT, our extended SwinBERT’s CIDEr score improves by 25.3% compared to without extension, while the extension of GL-RG exhibits an increase of 33.7%. These demonstrate the effectiveness of our utilization of user-specified targets. As these are intuitive extensions, we believe future dedicatedly designed methods could further boost the caption quality. (III) The performance of our extended methods still fall far behind their performance on the original video captioning tasks. For example, our best-performing extended SwinBERT on the subject-oriented MSVD, scored lower across all four performance metrics compared to the original SwinBERT in the conventional video captioning task, with a notable 35.2% decrease in BLEU@4. These indicate there is a large room for future exploration in this field.
Subject Encoder SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B (Faster-RCNN) 30.3 23.8 52.5 76.2 19.3 17.6 41.8 50.8
AR-B (CLIP V/L-14) 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN (Faster-RCNN) 32.7 24.1 53.6 85.2 19.8 17.6 42.5 52.2
HMN (CLIP V/L-14) 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG (Faster-RCNN) 37.8 27.5 56.5 95.9 20.8 18.2 42.0 53.3
GL-RG (CLIP V/L-14) 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert (Faster-RCNN) 35.6 25.9 55.9 98.1 21.3 19.1 43.9 58.0
SwinBert (CLIP V/L-14) 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 3: Results under different visual subject encoders.
Effect of Different Subject Feature
In the computer vision community, there are two popular object encoders: Faster-RCNN and VIT-CLIP. To investigate the effect of different subject features on performance, we conduct tests using Faster-RCNN and VIT(L-14)-based CLIP to extract object features, respectively. The results are presented in Table 3. It shows that CLIP V/L-14 leads to better results for all methods on the SO-MSVD dataset in terms of the main metric Cider. The improvements range from 0.8 to 10.7. On the SO-MSRVTT dataset, the performance of three out of four methods is improved. Overall, using CLIP-VIT features achieves better performance.
Effect of Different Manners of Utilizing Subject Features
In addition to concatenating subject features with video features, we have also explored using addition to impose the visual subjects on video features. The results are present in Table 4. As shown in Table 4, for methods based on LSTM (i.e., HMN and GL-RG), using concatenation significantly respectively enhances CIDEr performance 4.5% and 19.0% on the SO-MSVD dataset. In contrast, for Transformer-based methods, the difference between concatenation and addition is marginal.
The main reason can be attributed to the fact that in LSTM-based methods, the length of subject features has to be the same as the video feature for the addition operation, while it becomes more flexible in the concatenation operation. In contrast, in Transformer-based methods, as the concatenation is performed along the sequence dimension, the feature dimension has to be the same as other tokens.
Combination Operations SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B (Add) 33.0 25.1 53.6 88.2 19.4 17.7 40.5 49.6
AR-B (Concat) 33.5 25.3 53.8 86.9 19.6 17.8 40.6 50.2
HMN (Add) 31.9 23.9 53.0 84.3 19.3 17.3 41.2 50.3
HMN (Concat) 34.3 24.8 54.3 88.1 20.2 17.9 42.8 53.1
GL-RG (Add) 35.9 25.5 54.5 82.0 22.1 18.9 43.6 55.9
GL-RG (Concat) 37.2 27.8 56.9 97.6 22.2 18.9 43.9 58.6
SwinBert (Add) 36.6 27.2 56.5 97.2 21.7 18.9 44.5 59.0
SwinBert (Concat) 36.1 27.2 56.3 98.9 21.5 19.1 44.0 58.8
Table 4: Results under different combination operations based on CLIP subject features.
Effect of Different Dimension of Subject Features
Figure 7: Results (CIEDr score) on Subject-Oriented MSVD and MSRVTT under different subject feature dimensions.
For LSTM-based methods, the dimensionality of the subject features can significantly impact the performance of the generated captions, as illustrated in Figure 7. For SO-MSVD, lower dimensions of subject features tend to perform better, while for SO-MSRVTT, higher dimensions show improved performance. This is because the SO-MSRVTT dataset is larger and more complex, and higher feature dimensions help the model to better fit and capture the diversity and complexity of subject features.
Figure 8: Qualitative results on Subject-Oriented MSR-VTT.
Accuracy of Predicted Subject
In addition to using four conventional video description quality assessment metrics, we also compared the subject accuracy of the modified methods against the original methods on the test set of the new datasets. If the predicted subject is found within the set of ground truth subjects, the prediction is considered correct; otherwise, it is deemed incorrect.
The results, as shown in Table 5, indicate that our method achieves improvements over the baseline. Specifically, on the SO-MSRVTT dataset, SwinBERT and GL-RG achieved significant enhancements of approximately 12.8% and 11.8%, respectively. Meanwhile, on the SO-MSVD dataset, AR-B showed an improvement of around 9.8%.
@@ 6.2 Quantitative Results
GL-RG (Ours) 60.0% 57.4%
SwinBert 60.7% 48.8%
SwinBert (Ours) 63.9% 61.6%
Table 5: Subject accuracy of different methods on our Subject-Oriented MSVD and MSR-VTT benchmarks.
6.3 Qualitative Results
We show several qualitative results on our subject-oriented MSRVTT in Figure 8. Our improved method, compared to the four models that do not utilize subject areas, better focuses on and accurately describes the entities of our interest. For instance, the video in Figure 8 (a) shows three people conversing. It can be observed that the four original methods only described two of them. In contrast, our method generates accurate descriptions for the individually marked single, double, and triple-person groups in the video. Figure 8 (b) displays a group of people visiting a large boat. SwinBERT incorrectly identifies the boat in the video as “A large train”, while the other three methods describe the people. Our method can accurately describe both “A boat is being shown” and “a man is watching a boat”. In Figure 8 (c), the basic methods only describe people. However, our method not only describes “A woman is talking” and “A couple of people are walking down the street”, but also notes “A white building is being filmed”. In Figure 8 (d), our method is capable of describing “A machine” appearing in the video, an entity not mentioned by the original methods. The qualitative results further prove the effectiveness of our method and the significance of this task.
6.4 Future Work
The task of subject-oriented video caption generation is highly challenging. It requires the model to not only understand the video content but also to focus specifically on the behavior of entities that are of interest to the user. How to more effectively utilize the subject area is one of the possible research directions for this task. Future work may also explore large-scale vision-language models, such as GPT-4V.
7 Conclusion
In this paper, we introduce a new task, Subject-Oriented video captioning, which aims at describing the activities of entities in videos that are of interest to users. We build two datasets to support future research in this field and evaluate four state-of-the-art conventional video captioning methods as well as our extensions based on them. Experimental results show a large performance drop of existing methods on our new tasks and significant improvement of our extended models. Our extended methods can serve as strong baselines, but there is still a large room worth exploring for future works.

### v2

@@ V-A2 Baseline models
V-A3 Implementation Details
For the aforementioned four fundamental video captioning methods, we employ CLIP or Faster R-CNN as the subject encoder. Among these, we further extend the highest-performing SwinBERT.
In the subject-oriented sampling module, the number of clusters is $32$. In the subject-oriented encoding module, we set $5$ learnable prompt tokens. The batch size and learning rate are set to $16$ and $7.5\times 10^{-5}$ on SO-MSRVTT and SO-MSVD. We keep other parameters consistent with the original experimental settings of SwinBERT. The experiments are conducted on multiple Nvidia RTX 3090Ti GPUs and A6000 GPUs.
TABLE II: Results on our Subject-Oriented MSVD and MSR-VTT datasets. Rows with denote results on original MSVD and MSRVTT datasets. The best results are highlighted in bold.
@@ V-A3 Implementation Details
AR-B [58]_ori 2021 48.7 33.3 - 91.8 40.5 28.7 - 49.1
AR-B 2024 27.4 22.4 50.5 69.7 17.0 16.6 38.7 43.3
AR-B_ext 2024 31.6 24.4 53.2 78.2 20.6 18.5 42.9 55.5
HMN [12]_ori 2022 59.2 37.7 75.1 104.0 43.5 29.0 62.7 51.5
HMN 2024 29.8 22.9 51.2 77.7 17.1 16.5 39.1 44.9
HMN_ext 2024 34.3 24.8 54.3 88.1 20.1 17.9 42.8 53.1
GL-RG [31]_ori 2022 45.5 30.1 62.6 94.3 55.5 37.8 74.7 51.2
GL-RG 2024 34.5 26.1 54.4 86.7 18.3 17.3 39.1 43.8
GL-RG_ext 2024 36.2 27.8 56.9 96.6 22.1 18.8 43.7 56.4
SwinBert [11]_ori 2022 55.7 39.7 75.7 109.4 41.9 29.9 62.1 53.8
SwinBert 2024 33.4 25.7 54.2 89.6 18.2 17.4 40.4 46.9
SwinBert_ext 2024 35.8 27.1 56.3 97.7 21.5 18.9 44.3 57.1
SOVCNet (Ours) 2024 36.5 27.6 57.6 98.8 22.1 19.1 45.5 61.1
TABLE III: Subject accuracy of different methods on our Subject-Oriented MSVD and MSR-VTT benchmarks.
@@
GL-RG_ext 60.0% 57.4%
SwinBERT 60.9% 48.8%
SwinBERT_ext 63.9% 61.0%
SOVCNet 65.7% 63.7%
V-B Subject-Oriented Video Captioning Performance
To the best of our knowledge, no methods are doing precisely the same task as ours before. Therefore, we use the four extended methods mentioned in Section V-A2 for comparison. Table II shows the performance comparison between our method and other baseline models on the SO-MSVD and SO-MSRVTT test sets. For each extended baseline model, the first and second rows present their performance on the origin datasets and our reproduced datasets, respectively. The results show that: (I) The four state-of-the-art methods encounter a large performance drop on our new task compared to the performance on the original video captioning task. For example, on our subject-oriented MSVD, the CIDEr score of AR-B decreases by 31.7% relative to itself, while HMN reduces by 33.8% compared to its own performance. These indicate our subject-oriented video captioning task poses new challenges to existing methods. To succeed, they should have the ability to be aware of the specified target, and the ability to distinguish related activities. (II) Our extension of the four state-of-the-art methods brings large improvement. For example, on subject-oriented MSRVTT, our extended SwinBERT’s CIDEr score improves by 21.3% compared to without extension, while the extension of GL-RG exhibits an increase of 28.8%. These demonstrate the effectiveness of our utilization of user-specified targets. As these are intuitive extensions, we believe future dedicatedly designed methods could further boost the caption quality. (III) Our SOVCNet achieves the best performance across all metrics compared to the other four extended methods. Specifically, compared to the second-best SwinBert_ext, our SOVCNet gains 4.0 and 1.1 absolute improvements in terms of the main metric CIDEr on SO-MSRVTT and SO-MSVD, respectively. Meanwhile, we observe 0.6 and 0.7 absolute improvements in terms of another main metric BLEU@4. The large improvement demonstrates the effectiveness of our proposed SOVCNet.
In addition to using four conventional video description quality assessment metrics, we also compared the subject accuracy of SOVCNet against the original methods on the test set of the new datasets. If the predicted subject is found within the set of ground truth subjects, the prediction is considered correct; otherwise, it is deemed incorrect.
As shown in Table III, the accuracy of the four extended methods has also further improved compared to the original four methods. Additionally, our SOVCNet achieves the highest accuracy of 65.7% and 63.7% on SO-MSVD and SO-MSRVTT, leading by 1.8% and 2.7% absolute improvements over the second-best SwinBERT_ext. This shows that SOVCNet not only produces higher-quality captions but also more accurately predicts the subject specified by the user.
TABLE IV: Performance comparison of different modules in SOVCNet on SO-MSVD and SO-MSRVTT. The best results are highlighted in bold.
Model SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
SwinBERT_ext (Baseline) 35.8 27.1 56.3 97.7 21.5 18.9 44.3 57.1
SwinBERT_ext + SOS 36.5 27.1 56.4 98.3 21.7 19.1 44.4 59.2
SwinBERT_ext + SOE 35.2 26.9 57.5 96.6 20.6 19.2 43.4 51.9
SOVCNet (Ours) 36.5 27.6 57.6 98.8 22.1 19.1 45.5 61.1
TABLE V: A comparison of effect of different sampling strategies on SO-MSRVTT.
Sampling strategy B@4 M R C
Regular 20.6 19.2 43.4 51.9
Similarity 21.5 19.2 44.8 58.1
Adding-interval 21.9 19.3 45.0 59.2
Clustering 22.1 19.1 45.5 61.1
V-C Ablation Study
In this section, we present the ablation study results of our SOVCNet. We first ablate the effectiveness of our proposed SOS and SOE modules. Then, we demonstrate the impact of different SOS strategies and the number of SOE soft tokens. Finally, we show the effect of using different subject features in the generation phase.
Effectiveness of Proposed Modules
To evaluate the effectiveness of our proposed key modules on overall performance, we do ablation experiments on both SO-MSVD and SO-MSRVTT. The results are presented in Table IV. The SOS module makes the greatest contribution, increasing the CIDEr score to 59.2 from 57.1 on SO-MSRVTT, and to 98.3 from 97.7 on SO-MSVD. Our full model achieves the best performance, with the CIDEr score reaching 61.1 and 98.8 on SO-MSRVTT and SO-MSVD. However, performance drops when solely using SOE. This may be because we adjust the model’s frame encoding process without altering the frame sampling, resulting in a reduced ability to encode standard frames.
Effect of different sampling strategies
To study the effect of different sampling strategies on our method, we compare the performance of four different sampling strategies on SO-MSRVTT. The first row represents regular uniform sampling. The second row converts the similarity between all frames and the subject frames into probabilities for probabilistic sampling. The third row adds intervals on the basis of the second row to increase the diversity of information. The last row clusters video frames into k categories, and then probabilistically selects one frame from each category. The comparison results show that our cluster-based sampling method achieves the highest CIDEr score of 61.1, exceeding other methods by 1.9 to 9.2.
Figure 8: Qualitative results on Subject-Oriented MSR-VTT.
Figure 9: Comparison of number of Soft tokens on SO-MSRVTT.
TABLE VI: Results under different visual subject encoders.
Method SO-MSVD SO-MSRVTT
B@4 M R C B@4 M R C
AR-B_ext (Faster-RCNN) 30.3 23.8 52.5 76.2 19.3 17.6 41.8 50.8
AR-B_ext (CLIP V/L-14) 31.6 24.4 53.2 78.2 20.6 18.5 42.9 55.5
HMN_ext (Faster-RCNN) 32.7 24.1 53.6 85.2 19.8 17.6 42.5 52.2
HMN_ext (CLIP V/L-14) 34.3 24.8 54.3 88.1 20.1 17.9 42.8 53.1
GL-RG_ext (Faster-RCNN) 37.8 27.5 56.5 95.9 20.8 18.2 42.0 53.3
GL-RG_ext (CLIP V/L-14) 36.2 27.8 56.9 96.6 22.1 18.8 43.7 56.4
SwinBERT_ext (Faster-RCNN) 35.6 26.7 56.2 96.8 21.2 19.1 44.1 56.8
SwinBERT_ext (CLIP V/L-14) 35.8 27.1 56.3 97.7 21.5 18.9 44.3 57.1
SOVCNet (Faster-RCNN) 35.9 27.1 56.6 97.8 21.4 19.1 44.5 58.1
SOVCNet (CLIP V/L-14) 36.5 27.6 57.6 98.8 22.1 19.1 45.5 61.1
Effect of the number of soft tokens
We explore the effect of the number of soft tokens, and Figure 9 shows that 5 is the better choice than other numbers. In Figure 9, the CIDEr score initially drops as the number increases from 1 to 3, then reaches its highest point from 3 to 5, and subsequently falls again from 5 to 7. Consequently, we set 5 as the standard number of soft tokens.
Effect of different subject features
In Table VI, we compare the performance of our method using Faster-RCNN and VIT(L-14)-based CLIP to extract subject features, respectively. It shows that CLIP V/L-14 leads to better results for all methods on the SO-MSVD dataset in terms of the main metric CIDEr. The improvements range from 0.9 to 2.9. And the performance of the SO-MSRVTT dataset also significantly improves. For example, SOVCNet increases the CIDEr score on SO-MSRVTT from 58.1 to 61.1. Overall, using CLIP-VIT features achieves better performance.
V-D Qualitative Results
We show some qualitative results in Figure 8. Compared to the origin SwinBERT and extended SwinBERT, our SOVCNet better focuses on and accurately describes the entities of interest. For instance, the video in Figure 8 (a) shows a voice show. It can be observed that the origin SwinBERT only describes “a girl”. Compared to the extended SwinBERT, our method generates accurate descriptions for the girl, the judges, and one of the judges, as well as their actions. Figure 8 (b) displays two men talking about a boxing match. The extended SwinBERT incorrectly identified two speakers in a conversation as “a man is talking about a wrestling match”, while SOVCNet accurately describes it as “two men are talking”. In Figure 8 (c), the extended SwinBERT only roughly describes “a plate of food is being cooked in a restaurant”. in contrast, our improved method captures “a grill” in more detail. In Figure 8 (d), compared to Pred2 “a man is singing a song on stage” in extended SwinBERT, SOVCNet is capable of describing the “playing guitar” action appearing in the video. The qualitative results further prove the effectiveness of our method and the significance of this task.
VI Conclusion
In this paper, we propose a novel subject-oriented video captioning task that requires the model to generate subject-oriented video descriptions based on the subject specified by the user. To solve this task, we propose a corresponding method called SOVCNet. This approach begins by sampling frames related to the subject to minimize irrelevant information. Then it uses the frame containing the subject as hard prompts and plugs the soft prompts to enhance the model’s focus on the subject’s activities, aiding adaptation to the generative task. Additionally, we incorporate extra subject information to help the model generate accurate descriptions. To evaluate our system, we construct two new datasets using the MSVD and MSRVTT datasets. Qualitative and quantitative experimental results demonstrate the effectiveness of our method.